# Design of Bayesian A/B Tests Controlling False Discovery Rates and Power

### v4

@@ Abstract
Online controlled experiments (i.e., A/B tests) are a critical tool used by businesses with digital operations to optimize their products and services. These experiments routinely track information related to various business metrics, each of which summarizes a different aspect of how users interact with an online platform. Although multiple metrics are commonly tracked, this information is often not well utilized; multiple metrics are often aggregated into a single composite measure, losing valuable information, or strict family-wise error rate adjustments are imposed, leading to reduced power. In this paper, we propose an economical framework to design Bayesian A/B tests while controlling both power and the false discovery rate (FDR). Selecting optimal decision thresholds to control power and the FDR typically relies on intensive simulation at each sample size considered. Our framework efficiently recommends optimal sample sizes and decision thresholds for Bayesian A/B tests that satisfy criteria for the FDR and average power. Our approach is efficient because we leverage new theoretical results to obtain these recommendations using simulations conducted at only two sample sizes. Our methodology is illustrated using an example based on a real A/B test involving several metrics.
@@ 1 Introduction
For each metric $k=1,\dots,K$, A/B tests compare complementary hypotheses $H_{0,k}$ and $H_{1,k}$. A discovery is made with respect to the $k^{\text{th}}$ metric when the data support $H_{1,k}$. A/B tests with multiple metrics are often designed to control the family-wise error rate (FWER). However, strict control of the FWER often results in a loss of power. Since insufficient power in A/B tests is a common problem (Deng et al., 2013; Larsen et al., 2024; Deng et al., 2024), any design choices that further decrease power should be avoided. Thus, we control the false discovery rate (FDR): the expected proportion of discoveries that are false across a group of hypothesis tests (Benjamini and Hochberg, 1995). Whereas the FWER conservatively bounds the probability of making even one false discovery (i.e., incorrectly concluding that one $H_{1,k}$ is true), the FDR balances the prevalence of false and true positives across multiple comparisons. The development of design methods for A/B tests with FDR control is therefore crucial to enhance the average power of these tests. Average power is the expected proportion of the true $\{H_{1,k}\}_{k=1}^{K}$ that are correctly supported by the data.
@@
The most popular methods used to control the FDR are based on frequentist $p$-values (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001; Storey, 2002). Given the rising prominence of Bayesian approaches to A/B testing (see e.g., Deng (2015); Stevens and Hagar (2022); Deng et al. (2024)), this paper focuses on testing procedures facilitated via posterior probabilities. For the $k^{\text{th}}$ metric, the observed data provide evidence to support $H_{1,k}$ if the posterior probability $Pr(H_{1,k}~{}|~{}data)$ is greater than or equal to a decision threshold $\gamma_{k}\in[0.5,1)$. Given some data generation process, a sample size $n$ could be selected to ensure the average power to observe $Pr(H_{1,k}~{}|~{}data)\geq\gamma_{k}$ across all true hypotheses in $\{H_{1,k}\}_{k=1}^{K}$ is at least $1-\beta$. The decision thresholds $\boldsymbol{\gamma}=\{\gamma_{k}\}_{k=1}^{K}$ are chosen to bound the FDR by $q\in(0,1)$. The FDR control provided by this framework with posterior probabilities is compared with the FDR control prompted by popular frequentist methods in Section 2.
@@
To support the design of A/B tests with non-simplistic statistical models, sample sizes and optimal decision thresholds that control the operating characteristics of average power and the FDR can be found using intensive simulation (Wang and Gelfand, 2002). In general, one must simulate many samples of size $n$ according to some assumed data generation process to estimate the joint sampling distribution of $\{Pr(H_{1,k}~{}|~{}data)\}_{k=1}^{K}$. This estimate of the sampling distribution of posterior probabilities determines whether the investigated sample size and decision thresholds satisfy criteria for the operating characteristics. This computationally intensive process is repeated until a suitable $(n,\boldsymbol{\gamma})$ combination is found. An economical framework to determine optimal sample sizes and decision thresholds that satisfy criteria for the FDR and average power would expedite the design of Bayesian A/B tests with multiple metrics, making them more accessible to practitioners.
@@
For a single test, Hagar and Stevens (2025) recently put forward an economical method to assess operating characteristics throughout the sample size space using estimates of the sampling distribution of posterior probabilities at only two sample sizes. In this work, we build upon their approach to more efficiently design A/B tests involving a host of metrics controlling the FDR and average power. To assess the operating characteristics of multiple tests, we must account for the dependence structure in the joint sampling distribution of posterior probabilities across hypotheses. The method from Hagar and Stevens (2025) was predicated on a theoretical proxy to the (univariate) sampling distribution of posterior probabilities for a single hypothesis. As such, new theory must be developed to accommodate arbitrary dependence structures in the joint sampling distribution.
@@
The remainder of this article is structured as follows. We introduce our framework for inference and the criteria under which the FDR and average power are controlled in Section 2. In Section 3, we prove novel theoretical results about a proxy to the joint sampling distribution of posterior probabilities. We develop a method in Section 4 that adapts these results to determine which $(n,\boldsymbol{\gamma})$ combination minimizes the sample size $n$ while satisfying criteria for average power and the FDR. This method efficiently considers a range of sample sizes using estimates of the sampling distributions of posterior probabilities at only two values of $n$. In Section 5, we illustrate the use of our methodology with an example from a real A/B test with multiple binary metrics that have a complicated dependence structure. We conclude with a summary and discuss extensions to this work in Section 6.
@@ 2.1 Problem Set-Up
The targets of inference $\boldsymbol{\theta}$ for the host of metrics are specified as a function $g(\cdot)$ of the model parameters: $\boldsymbol{\theta}=g(\boldsymbol{\eta})$. The components in $\boldsymbol{\theta}=\{\theta_{k}\}_{k=1}^{K}$ can be expressed as $\theta_{k}=g_{k}(\boldsymbol{\eta})$. For the $k^{\text{th}}$ metric, we consider interval hypotheses of the form $H_{1,k}:\theta_{k}\in(\delta_{L,k},\delta_{U,k})$, where $-\infty\leq\delta_{L,k}<\delta_{U,k}\leq\infty$. The corresponding complementary hypothesis is $H_{0,k}:\theta_{k}\notin(\delta_{L,k},\delta_{U,k})$. This general notation for the interval $(\delta_{L,k},\delta_{U,k})$ accommodates a broad suite of hypothesis tests based on superiority and practical equivalence (Stevens and Hagar, 2022). We jointly refer to the interval endpoints across all hypotheses as $\boldsymbol{\delta}_{L}=\{\delta_{L,k}\}_{k=1}^{K}$ and $\boldsymbol{\delta}_{U}=\{\delta_{U,k}\}_{k=1}^{K}$.
@@
Algorithm 1 details a simulation-driven procedure to classify hypotheses as true or false for a particular $(n,\boldsymbol{\gamma})$ combination and set of design inputs. These classifications based on posterior probabilities are used to estimate the FDR and average power. We now discuss several design inputs for this algorithm. The prior $p(\boldsymbol{\eta})$ is used to define the posterior of $\boldsymbol{\theta}$. We also specify $\Psi$, a probability model that characterizes how $\boldsymbol{\eta}$ values are drawn in each simulation repetition $r=1,\dots,m$. The model $\Psi$ defines scenarios with effect sizes that align with the experiment’s objectives. It differs from the prior $p(\boldsymbol{\eta})$ that characterizes all available pre-experimental knowledge regardless of whether it aligns with the goals of the A/B tests. As illustrated in Section 5, defining a model $\Psi$ that is distinct from $p(\boldsymbol{\eta})$ allows practitioners to incorporate uncertainty about how many and which hypotheses are false in a straightforward manner. Note that the FDR is trivially 0 if all hypotheses $\{H_{1,k}\}_{k=1}^{K}$ are true, and average power is trivially 0 when all $\{H_{1,k}\}_{k=1}^{K}$ are false. In this paper, we therefore focus on data generation processes where each $\boldsymbol{\eta}_{r}\sim\Psi$ is such that some – but not all – of $\{H_{1,k}\}_{k=1}^{K}$ are false. We use this assumption about $\{H_{1,k}\}_{k=1}^{K}$ for design purposes to ensure the FDR and average power calculations are non-trivial in all simulation repetitions; there are no restrictions on how many hypotheses are false when using posterior probabilities to assess a collection of A/B tests. We note that one could also use our design framework with a model $\Psi$ such that all $\{H_{1,k}\}_{k=1}^{K}$ are true or false in some simulation repetitions $r=1,\dots,m$.
@@
1:procedure Classify($f(\cdot)$, $g(\cdot)$, $\boldsymbol{\delta}_{L}$, $\boldsymbol{\delta}_{U}$, $n$, $c$, $\boldsymbol{\gamma}$, $K$, $p(\boldsymbol{\eta})$, $m$, $\Psi$)
@@
11:              else
12:                  $t_{r}\leftarrow t_{r}+1$
13:                  if $\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})\geq\gamma_{k}$ then
14:                       $s_{r}\leftarrow s_{r}+1$
15:     return $\{v_{r}\}_{r=1}^{m}$, $\{s_{r}\}_{r=1}^{m}$, and $\{t_{r}\}_{r=1}^{m}$
Line 4 of Algorithm 1 generates a sample $\boldsymbol{y}^{{}_{(n)}}_{r}$ in each simulation repetition given the value for $\boldsymbol{\eta}_{r}$ drawn from $\Psi$ in Line 3. The collection of $\{\{\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ values from Line 7 across all hypotheses and simulation repetitions estimates the joint sampling distribution of posterior probabilities under the model $\Psi$. The operating characteristics of the FDR and average power can respectively be estimated as
@@
where $v_{r}$, $s_{r}$, and $t_{r}$ are the number of false discoveries, true discoveries, and true hypotheses for simulation repetition $r$ defined in Algorithm 1. The denominators in (1) avoid division by 0, aligning with the definition of the FDR from Benjamini and Hochberg (1995). Based on Algorithm 1, an $(n,\boldsymbol{\gamma})$ combination is suitable if the FDR estimate in (1) is at most $q$ and the average power estimate in (1) is at least $1-\beta$.
@@
The optimal decision thresholds $\boldsymbol{\gamma}$ that bound the estimated FDR in (1) by $q$ can be found using a defined optimization scheme. We illustrate the use of several constrained optimize schemes that select $\boldsymbol{\gamma}$ to maximize average power in Section 5. As $n$ increases, so do the posterior probabilities in the components of the marginal sampling distributions where $H_{1,k}$ is true, and it becomes easier to distinguish between the true and false hypotheses. The optimal $\boldsymbol{\gamma}$ values for a given optimization scheme typically decrease alongside $n$; however, this decreasing trend is attenuated as $n\rightarrow\infty$ since the components of the marginal sampling distributions of posterior probabilities where $H_{1,k}$ is true converge to 1. Even as the optimal $\boldsymbol{\gamma}$ values change, average power increases to 1 as $n$ increases. The optimal choices for $n$ and $\boldsymbol{\gamma}$ are therefore intrinsically linked when designing A/B tests based on the FDR and average power.
@@
These choices for $n$ and $\boldsymbol{\gamma}$ are informed by exploring the joint sampling distribution of posterior probabilities under $\Psi$ to estimate the operating characteristics in (1). While we can examine multiple $\boldsymbol{\gamma}$ values using the same estimate of the sampling distribution for a given sample size $n$, the procedure in Algorithm 1 requires independent implementation for each $n$ value we consider. This process is often computationally intensive, but we could reduce the computational burden by repurposing the results from Algorithm 1 for previously considered sample sizes to estimate operating characteristics for new $n$ values. This process would allow us to explore new $(n,\boldsymbol{\gamma})$ combinations without conducting additional simulations. We begin the development of such a method for the design of A/B tests in Section 3. But first, we compare the FDR control provided by our design framework with the FDR control prompted by standard methods.
2.2 False Discovery Rate Control
To compare and contrast how the FDR is controlled in our design framework and under
alternative approaches, we let $V$ and $S$ be random variables that denote the number of true and false discoveries, respectively. Our design method controls the FDR such that
$\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\lim_{n\rightarrow\infty}\mathbb{E}\left[\dfrac{V}{V+S}\right]\right\}\leq q,$ (2)
where $0\div 0$ is defined to be 0. We emphasize that the FDR is controlled in expectation under the assumption that $\boldsymbol{\eta}\sim\Psi$, and we discuss the implications when this assumption does not hold true in Section 5. Our FDR control is also defined with respect to the model $f(\boldsymbol{y};\boldsymbol{\eta})$ for the data that practitioners must select. The criterion in (2) holds true as $n\rightarrow\infty$ because our method to explore the sample size space leverages asymptotic theory. Nevertheless, our method performs well with finite $n$ since we estimate sampling distributions with simulation.
where $0\div 0$ is again defined to be 0. The criterion in (3) holds for an arbitrary value for the parameter $\boldsymbol{\eta}$ that indexes the data generation process. Standard methods for FDR control assume that $p$-values are uniformly distributed under each $H_{0,k}$. This uniformity is often justified asymptotically using the central limit theorem, so standard methods for FDR control do not require practitioners to choose a model $f(\boldsymbol{y};\boldsymbol{\eta})$ for the data. However, the FDR control in Benjamini and Hochberg (1995) is only attained for certain dependence structures, including settings where all the hypotheses are independent. Benjamini and Yekutieli (2001) developed a method that attains the criterion in (3) for an arbitrary dependence structure, but simulation-based procedures are recommended for optimal control of the FDR for a specific dependence structure (Yekutieli and Benjamini, 1999). Ultimately, our design framework requires users to make more choices about the data generation process and model for the data; the trade-off of using our methods is that the recommended sample sizes and decision thresholds are optimized when these choices are reasonable.

### v5

@@ Abstract
Businesses frequently run online controlled experiments (i.e., A/B tests) to learn about the effect of an intervention on multiple business metrics. To account for multiple hypothesis testing, multiple metrics are commonly aggregated into a single composite measure, losing valuable information, or strict family-wise error rate adjustments are imposed, leading to reduced power. In this paper, we propose an economical framework to design Bayesian A/B tests while controlling both power and the false discovery rate (FDR). Selecting optimal decision thresholds to control power and the FDR typically relies on intensive simulation at each sample size considered. Our framework efficiently recommends optimal sample sizes and decision thresholds for Bayesian A/B tests that satisfy criteria for the FDR and average power. Our approach is efficient because we leverage new theoretical results to obtain these recommendations using simulations conducted at only two sample sizes. Our methodology is illustrated using an example based on a real A/B test involving several metrics.
@@ 1 Introduction
For each metric $k=1,\dots,K$, A/B tests compare complementary hypotheses $H_{0,k}$ and $H_{1,k}$. A discovery is made with respect to the $k^{\text{th}}$ metric when the data support $H_{1,k}$. A/B tests with multiple metrics are often designed to control the family-wise error rate (FWER). However, strict control of the FWER often results in a loss of power. Since insufficient power in A/B tests is a common problem (Deng et al., 2013; Larsen et al., 2024; Deng et al., 2024), any design choices that further decrease power should be avoided. Thus, we control the false discovery rate (FDR): the expected proportion of discoveries that are false across a group of hypothesis tests. Whereas the FWER conservatively bounds the probability of making even one false discovery (i.e., incorrectly concluding that one $H_{1,k}$ is true), the FDR balances the prevalence of false and true positives across multiple comparisons. The development of design methods for A/B tests with FDR control is therefore crucial to enhance the average power of these tests. Average power is the expected proportion of the true $\{H_{1,k}\}_{k=1}^{K}$ that are correctly supported by the data.
@@
The most popular methods used to control the FDR are based on frequentist $p$-values (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001; Storey, 2002). Given the rising prominence of Bayesian approaches to A/B testing (see e.g., Deng (2015); Stevens and Hagar (2022); Deng et al. (2024)), this paper focuses on testing procedures facilitated via posterior probabilities. For the $k^{\text{th}}$ metric, the observed data provide evidence to support $H_{1,k}$ if the posterior probability $Pr(H_{1,k}~|~data)$ is greater than or equal to a decision threshold $\gamma_{k}\in[0.5,1)$. Given some data generation process, a sample size $n$ could be selected to ensure the average power to observe $Pr(H_{1,k}~|~data)\geq\gamma_{k}$ across all true hypotheses in $\{H_{1,k}\}_{k=1}^{K}$ is at least $1-\beta$. The decision thresholds $\boldsymbol{\gamma}=\{\gamma_{k}\}_{k=1}^{K}$ are chosen to bound the FDR by $q\in(0,1)$. The Bayesian FDR control provided by this framework with posterior probabilities is defined and compared with the FDR control prompted by popular frequentist methods in Section 2.
@@
To support the design of A/B tests with non-simplistic statistical models, sample sizes and optimal decision thresholds that control the operating characteristics of average power and the Bayesian FDR can be found using intensive simulation (Wang and Gelfand, 2002). In general, one must simulate many samples of size $n$ according to some assumed data generation process to estimate the joint sampling distribution of $\{Pr(H_{1,k}~|~data)\}_{k=1}^{K}$. This estimate of the sampling distribution of posterior probabilities determines whether the investigated sample size and decision thresholds satisfy criteria for the operating characteristics. This computationally intensive process is repeated until a suitable $(n,\boldsymbol{\gamma})$ combination is found. An economical framework to determine optimal sample sizes and decision thresholds that satisfy criteria for the Bayesian FDR and average power would expedite the design of Bayesian A/B tests with multiple metrics, making them more accessible to practitioners.
@@
For a single test, Hagar and Stevens (2025) recently put forward an economical method to assess operating characteristics throughout the sample size space using estimates of the sampling distribution of posterior probabilities at only two sample sizes. In this work, we build upon their approach to more efficiently design A/B tests involving a host of metrics controlling the Bayesian FDR and average power. To assess the operating characteristics of multiple tests, we must account for the dependence structure in the joint sampling distribution of posterior probabilities across hypotheses. The method from Hagar and Stevens (2025) was predicated on a theoretical proxy to the (univariate) sampling distribution of posterior probabilities for a single hypothesis. As such, new theory must be developed to accommodate arbitrary dependence structures in the joint sampling distribution.
@@
The remainder of this article is structured as follows. We introduce our framework for inference and the criteria under which the Bayesian FDR and average power are controlled in Section 2. In Section 3, we prove novel theoretical results about a proxy to the joint sampling distribution of posterior probabilities. We develop a method in Section 4 that adapts these results to determine which $(n,\boldsymbol{\gamma})$ combination minimizes the sample size $n$ while satisfying criteria for average power and the Bayesian FDR. This method efficiently considers a range of sample sizes using estimates of the sampling distributions of posterior probabilities at only two values of $n$. In Section 5, we illustrate the use of our methodology with an example from a real A/B test with multiple binary metrics that have a complicated dependence structure. We conclude with a summary and discuss extensions to this work in Section 6.
@@ 2.1 Problem Set-Up
The targets of inference $\boldsymbol{\theta}$ for the host of metrics are specified as a function $g(\cdot)$ of the model parameters: $\boldsymbol{\theta}=g(\boldsymbol{\eta})$. The components in $\boldsymbol{\theta}=\{\theta_{k}\}_{k=1}^{K}$ can be expressed as $\theta_{k}=g_{k}(\boldsymbol{\eta})$. For the $k^{\text{th}}$ metric, we consider interval hypotheses of the form $H_{1,k}:\theta_{k}\in(\delta_{L,k},\delta_{U,k})$, where $\infty\leq\delta_{L,k}<\delta_{U,k}\leq\infty$. The corresponding complementary hypothesis is $H_{0,k}:\theta_{k}\notin(\delta_{L,k},\delta_{U,k})$. This general notation for the interval $(\delta_{L,k},\delta_{U,k})$ accommodates a broad suite of hypothesis tests based on superiority and practical equivalence (Stevens and Hagar, 2022). We jointly refer to the interval endpoints across all hypotheses as $\boldsymbol{\delta}_{L}=\{\delta_{L,k}\}_{k=1}^{K}$ and $\boldsymbol{\delta}_{U}=\{\delta_{U,k}\}_{k=1}^{K}$.
@@
We specify a probability model $\Psi$ that characterizes which $\boldsymbol{\eta}$ values define the data generation process. The model $\Psi$ defines scenarios with effect sizes that align with the experiment’s objectives. It differs from the prior $p(\boldsymbol{\eta})$ that defines the posterior of $\boldsymbol{\theta}$ and characterizes all available pre-experimental knowledge regardless of whether it aligns with the goals of the A/B tests. As illustrated in Section 5, defining a model $\Psi$ that is distinct from $p(\boldsymbol{\eta})$ allows practitioners to incorporate uncertainty about how many and which hypotheses are false in a straightforward manner.
To compare and contrast how the FDR is controlled in our design framework and under alternative approaches, we let $V$ and $S$ be random variables that denote the number of false and true discoveries, respectively. Our design method controls the FDR such that
$\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\mathbb{E}\left[\dfrac{V}{V+S}\right]\right\}\leq q,$ (1)
where $0\div 0$ is defined to be 0. In this paper, we refer to the FDR on the left side of the inequality in (1) as the Bayesian FDR. This Bayesian FDR is controlled in expectation under the assumption that $\boldsymbol{\eta}\sim\Psi$, and we discuss the implications when this assumption does not hold true in Section 5. Our Bayesian FDR control is also defined with respect to the model $f(\boldsymbol{y};\boldsymbol{\eta})$ for the data that practitioners must select.
where $0\div 0$ is again defined to be 0. The criterion in (2) holds for an arbitrary value for the parameter $\boldsymbol{\eta}$ that indexes the data generation process. Standard methods for FDR control assume that $p$-values are uniformly distributed under each $H_{0,k}$. This uniformity is often justified asymptotically using the central limit theorem, so standard methods for FDR control do not require practitioners to choose a model $f(\boldsymbol{y};\boldsymbol{\eta})$ for the data. However, the FDR control in Benjamini and Hochberg (1995) is only attained for certain dependence structures, including settings where all the hypotheses are independent. Benjamini and Yekutieli (2001) developed a method that attains the criterion in (2) for an arbitrary dependence structure, but simulation-based procedures are recommended for optimal control of the FDR for a specific dependence structure (Yekutieli and Benjamini, 1999). Ultimately, our design framework with Bayesian FDR control requires users to make more choices about the data generation process and model for the data; the trade-off of using our methods is that the recommended sample sizes and decision thresholds are optimized when these choices are reasonable.
To define our notion of average power, we let $T$ be a random variable that denotes the number of true hypotheses. Our design method controls average power across all true hypotheses in $\{H_{1,k}\}_{k=1}^{K}$ such that
$\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\mathbb{E}\left[\dfrac{S}{T}\right]\right\}\geq 1-\beta,$ (3)
where $0\div 0$ is defined to be 0. Standard design methods also control power with respect to some data generation process $\Psi$. The goal of our proposed method is to recommend a suitable $(n,\boldsymbol{\gamma})$ combination such that the criteria for the Bayesian FDR in (1) and average power in (3) are approximately attained. We formally describe how our method attains these criteria as the recommended value of $n\rightarrow\infty$ in Section 4. In the next subsection, we explain why standard methods to find suitable $(n,\boldsymbol{\gamma})$ combinations are very computationally intensive and how we improve upon standard design procedures.
2.2 A Computational Bottleneck
Algorithm 1 details a standard, simulation-driven procedure to classify hypotheses as true or false for a particular $(n,\boldsymbol{\gamma})$ combination and set of design inputs. These classifications based on posterior probabilities are used to estimate the Bayesian FDR and average power. Before explaining why design methods that repeatedly implement Algorithm 1 are computationally intensive, we elaborate on the choice of the model $\Psi$ that is input into this algorithm. Note that the Bayesian FDR is trivially 0 if all hypotheses $\{H_{1,k}\}_{k=1}^{K}$ are true, and average power is trivially 0 when all $\{H_{1,k}\}_{k=1}^{K}$ are false. In this paper, we therefore focus on data generation processes where each $\boldsymbol{\eta}_{r}\sim\Psi$ is such that some – but not all – of $\{H_{1,k}\}_{k=1}^{K}$ are false. We use this assumption about $\{H_{1,k}\}_{k=1}^{K}$ for design purposes to ensure the Bayesian FDR and average power calculations are non-trivial in all simulation repetitions; there are no restrictions on how many hypotheses are false when using posterior probabilities to assess a collection of A/B tests. We note that one could also use our design framework with a model $\Psi$ such that all $\{H_{1,k}\}_{k=1}^{K}$ are true or false in some simulation repetitions $r=1,\dots,m$.
@@ 2.2 A Computational Bottleneck
1:procedure Classify(f()f($\cdot$), $g(\cdot)$, $\boldsymbol{\delta}_{L}$, $\boldsymbol{\delta}_{U}$, $n$, $c$, $\boldsymbol{\gamma}$, $K$, $p(\boldsymbol{\eta})$, $m$, $\Psi$)
@@
11:      end if
12:     else
13:      $t_{r}\leftarrow t_{r}+1$
14:      if $\widehat{Pr}(H_{1,k}~|~\boldsymbol{y}^{{}_{(n)}}_{r})\geq\gamma_{k}$ then
15:        $s_{r}\leftarrow s_{r}+1$
16:      end if
17:     end if
18:   end for
19:  end for
20:  return $\{v_{r}\}_{r=1}^{m}$, $\{s_{r}\}_{r=1}^{m}$, and $\{t_{r}\}_{r=1}^{m}$
21:end procedure
Line 4 of Algorithm 1 generates a sample $\boldsymbol{y}^{{}_{(n)}}_{r}$ in each simulation repetition given the value for $\boldsymbol{\eta}_{r}$ drawn from $\Psi$ in Line 3. The collection of $\{\{\widehat{Pr}(H_{1,k}~|~\boldsymbol{y}^{{}_{(n)}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ values from Line 7 across all hypotheses and simulation repetitions estimates the joint sampling distribution of posterior probabilities under the model $\Psi$. The operating characteristics of the Bayesian FDR defined in (1) and average power defined in (3) can respectively be estimated as
@@
where $v_{r}$, $s_{r}$, and $t_{r}$ are the number of false discoveries, true discoveries, and true hypotheses for simulation repetition $r$ defined in Algorithm 1. The denominators in (4) avoid division by 0. Based on Algorithm 1, an $(n,\boldsymbol{\gamma})$ combination is suitable if the Bayesian FDR estimate in (4) is at most $q$ and the average power estimate in (4) is at least $1-\beta$.
@@
The optimal decision thresholds $\boldsymbol{\gamma}$ that bound the estimated Bayesian FDR in (4) by $q$ can be found using a defined optimization scheme. We illustrate the use of several constrained optimization schemes that select $\boldsymbol{\gamma}$ to maximize average power in Section 5. As $n$ increases, so do the posterior probabilities in the components of the marginal sampling distributions where $H_{1,k}$ is true, and it becomes easier to distinguish between the true and false hypotheses. The optimal $\boldsymbol{\gamma}$ values for a given optimization scheme typically decrease alongside $n$; however, this decreasing trend is attenuated as $n\rightarrow\infty$ since the components of the marginal sampling distributions of posterior probabilities where $H_{1,k}$ is true converge to 1. Even as the optimal $\boldsymbol{\gamma}$ values change, the estimated average power in (4) increases to 1 as $n$ increases. The optimal choices for $n$ and $\boldsymbol{\gamma}$ are therefore intrinsically linked when designing A/B tests based on the Bayesian FDR and average power.
@@
These choices for $n$ and $\boldsymbol{\gamma}$ are informed by exploring the joint sampling distribution of posterior probabilities under $\Psi$ to estimate the operating characteristics in (4). While we can examine multiple $\boldsymbol{\gamma}$ values using the same estimate of the sampling distribution for a given sample size $n$, the procedure in Algorithm 1 requires independent implementation for each $n$ value we consider. This process is often computationally intensive, but we could reduce the computational burden by repurposing the results from Algorithm 1 for previously considered sample sizes to estimate operating characteristics for new $n$ values. This process would allow us to explore new $(n,\boldsymbol{\gamma})$ combinations without conducting additional simulations. We begin the development of such a method for the design of A/B tests in Section 3.